# Electrically Tunable and Enhanced Nonlinearity of Moiré Exciton-Polaritons in Transition Metal Dichalcogenide Bilayers

### v1

@@ Abstract
$\Delta^{\bar{\beta}\bar{\alpha}}_{\sigma}=\sum_{\nu\nu^{\prime}}\sum_{\tilde{\nu}\tilde{\nu}^{\prime}}$ $\Big{[}\bar{C}^{\bar{\beta}}_{\nu^{\prime}\sigma}\bar{C}^{\bar{\alpha}}_{\tilde{\nu}^{\prime}\sigma}+\bar{C}^{\bar{\beta}}_{\tilde{\nu}^{\prime}\sigma}\bar{C}^{\bar{\alpha}}_{\nu^{\prime}\sigma}\Big{]}C^{\bar{\beta}}_{\nu\sigma}C^{\bar{\alpha}}_{\tilde{\nu}\sigma}$
$\Big{[}U_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}-V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}\Big{]}\delta_{\tilde{\bm{g}}^{\prime}+\bm{g}^{\prime},\tilde{\bm{g}}+\bm{g}}$ (9)
@@
$V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}=$ $\sum_{\mathbf{k}\tilde{\mathbf{k}}\mathbf{q}}\Gamma_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}(\mathbf{k}\tilde{\mathbf{k}},\mathbf{q})\delta_{\tilde{\ell}_{c}\ell_{c}^{\prime}}\delta_{\tilde{\ell}_{v}\tilde{\ell}_{v}^{\prime}}\delta_{\ell_{c}\tilde{\ell}_{c}^{\prime}}\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\mathbf{q},\tilde{\bm{g}}^{\prime}-\bm{g}+\mathbf{k}-\tilde{\mathbf{k}}},$
@@
Moiré exciton polariton.—Embedding the moiré bilayer into an optical microcavity, excitons can couple strongly to cavity photons, forming polaritonic states. This is an essential mechanism for studying optical nonlinearity [53, 54]. Here, dipolar polaritons in GaAs double quantum wells serve as an inspiration for studying pronounced nonlinear effects [55, 56, 57, 58]. To study moiré polaritons, we introduce light-matter coupling as an additional term in the system Hamiltonian corresponding to $\mathcal{H}_{\mathrm{sc}}=\sum_{\ell}d^{\ell}_{cv}\sum_{\mathbf{k}\mathbf{Q}}c^{\dagger}_{\mathbf{Q}}a^{\dagger}_{\ell\mathbf{k}}b_{\ell\mathbf{k}+\mathbf{Q}}+h.c.=\sum_{\bar{\alpha}\bar{\mathbf{Q}}}g^{\bar{\alpha}}_{c}(\bar{\mathbf{Q}})c^{\dagger}_{\bar{\mathbf{Q}}}\mathcal{X}^{\bar{\alpha}}(\bar{\mathbf{Q}})+h.c.$. Here $c^{\dagger}_{\mathbf{Q}}$ is the photonic field operator, and $d_{cv}^{\ell}$ is the interband transition matrix element in $\ell$-th layer. In the equation, this coupling term is written in the hybridized excitonic basis with coupling constant $g^{\bar{\alpha}}_{c}(\bar{\mathbf{Q}})=\sum_{\ell\alpha\mathbf{k}}d_{cv}^{\ell}\phi^{\ell\ell}_{\alpha}(\mathbf{k})\bar{C}^{\bar{\alpha},\ell\ell}_{\alpha\sigma}(\bar{\mathbf{Q}})$.
@@
In our analysis we focus on the nonlinear response and X-X interaction strength shown in Fig. 2b. In the plot we keep only the dominant contributions from Eq. (9) with $\bm{g}^{\prime}=\bm{g}$ and $\tilde{\bm{g}}^{\prime}=\tilde{\bm{g}}$ for simplicity, since the Coulomb scattering processes with nonzero momentum transfer between the excitons are rapidly suppressed as $\theta\gtrsim 0.5^{\circ}$. First of all, we find an enhancement in the nonlinear interaction at the small-$\theta$ regime indicating the strong influence of the band hybridization due to the Umklapp processes. We find strong attractive nonlinearity for hIX while the nonlinear interaction for hX remains repulsive. Intriguingly, these results correlate with measurements in MoS2 bilayers, where redshifts were observed [30]. The emergence of attractive nonlinearity may be understood as a manifestation of strong anisotropic nature of 2D bilayers. This leads to a weaker interlayer e-h interaction as compared to the intralayer e-e and h-h interaction (Fig. 2b, inset). Hence, the total nonlinearity from each of these interacting channels does not cancel leading to stronger attractive nonlinearity. This is different from the monolayer exciton where large cancellations [61, 62, 63] between these channels result in a weaker repulsive nonlinearity ($\sim 0.5$ nm2eV). This anisotropic electronic property is unique for 2D materials that enable the realization of attractive interaction between dipolar excitons. Furthermore, in our calculation, we find weak repulsive direct interaction for hIX in a homobilayer with $\theta\approx 0$ (see SM[52]). This may be expected since a large electrical dipole moment is forbidden due to the approximate inversion symmetry. As a result, the dominant contribution to $g_{X}$ in Fig. 2b comes from the (attractive) exchange interactions. However, this can change for heterobilayers or electrically-biased samples.
@@ Appendix C Momentum integration in direct and exchange interactions
$M^{\vec{r},\vec{t}}_{\vec{m}}(\vec{\lambda})=\Big{[}\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}\frac{(-1)^{r_{i}}m_{i}!(2\lambda_{i})^{m_{i}-2r_{i}}}{r_{i}!t_{i}!(m_{i}-2r_{i}-t_{i})!}\Big{(}\frac{2}{\vec{\lambda}^{2}}\Big{)}^{(m_{i}-2r_{i}-t_{i}+1)/2}\Big{]}\frac{1}{2}\Gamma\Big{(}\frac{\sum_{i=1}^{4}(m_{i}-2r_{i}-t_{i})+1}{2}\Big{)}.$ (38)
@@ Appendix D MoS2 homobilayer with $\theta\approx 0$
In this section, we calculate the spectrum of the MoS2 homobilayer with $\theta\approx 0$ in Fig.S1. In contrast to the $\theta\approx 60^{\circ}$ case, the attractive interaction in this stacking is stronger because the inversion symmetry forbid large electric dipole moment in a hybridized exciton.
Figure S1: (a) Moiré exciton spectrum with $\theta\approx 0$. (b) the nonlinearity of moiré exciton.

### v2

@@ Abstract
$\Delta^{\bar{\beta}\bar{\alpha}}_{\sigma}=\sum_{\nu\nu^{\prime}}\sum_{\tilde{\nu}\tilde{\nu}^{\prime}}$ $\bar{C}^{\bar{\beta}}_{\nu^{\prime}\sigma}\bar{C}^{\bar{\alpha}}_{\tilde{\nu}^{\prime}\sigma}C^{\bar{\beta}}_{\nu\sigma}C^{\bar{\alpha}}_{\tilde{\nu}\sigma}\delta_{\tilde{\bm{g}}^{\prime}+\bm{g}^{\prime},\tilde{\bm{g}}+\bm{g}}$
$\Big{[}U_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}+U_{\nu\tilde{\nu}}^{\tilde{\nu}^{\prime}\nu^{\prime}}-V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}-V_{\nu\tilde{\nu}}^{\tilde{\nu}^{\prime}\nu^{\prime}}\Big{]}$ (9)
@@
$V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}=$ $\sum_{\mathbf{k}\tilde{\mathbf{k}}\mathbf{q}}\Gamma_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}(\mathbf{k}\tilde{\mathbf{k}},\mathbf{q})\delta_{\tilde{\ell}_{c}\ell_{c}^{\prime}}\delta_{\tilde{\ell}_{v}\tilde{\ell}_{v}^{\prime}}\delta_{\ell_{c}\tilde{\ell}_{c}^{\prime}}\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\mathbf{q},\bm{g}^{\prime}-\bm{g}+\mathbf{k}-\tilde{\mathbf{k}}},$
@@
Moiré exciton polariton.—Embedding the moiré bilayer into an optical microcavity, excitons can couple strongly to cavity photons, forming polaritonic states. This is an essential mechanism for studying optical nonlinearity [53, 54]. Here, dipolar polaritons in GaAs double quantum wells serve as an inspiration for studying pronounced nonlinear effects [55, 56, 57, 58]. To study moiré polaritons, we introduce light-matter coupling as an additional term in the system Hamiltonian corresponding to $\mathcal{H}_{\mathrm{sc}}=\sum_{\ell}d^{\ell}_{cv}\sum_{\mathbf{k}\mathbf{Q}}c^{\dagger}_{\mathbf{Q}}a^{\dagger}_{\ell\mathbf{k}}b_{\ell\mathbf{k}+\mathbf{Q}}+h.c.=\sum_{\bar{\alpha}\bar{\mathbf{Q}}}g^{\bar{\alpha}}_{c}(\bar{\mathbf{Q}})c^{\dagger}_{\bar{\mathbf{Q}}}\mathcal{X}^{\bar{\alpha}}(\bar{\mathbf{Q}})+h.c.$ Here $c^{\dagger}_{\mathbf{Q}}$ is the photonic field operator, and $d_{cv}^{\ell}$ is the interband transition matrix element in $\ell$-th layer. In the equation, this coupling term is written in the hybridized excitonic basis with coupling constant $g^{\bar{\alpha}}_{c}(\bar{\mathbf{Q}})=\sum_{\ell\alpha\mathbf{k}}d_{cv}^{\ell}\phi^{\ell\ell}_{\alpha}(\mathbf{k})\bar{C}^{\bar{\alpha},\ell\ell}_{\alpha\sigma}(\bar{\mathbf{Q}})$.
@@
In our analysis we focus on the nonlinear response and X-X interaction strength shown in Fig. 2b. In the plot we keep only the dominant contributions from Eq. (9) with $\bm{g}^{\prime}=\bm{g}$ for simplicity, since the Coulomb scattering processes with nonzero momentum transfer between the excitons are rapidly suppressed as $\theta\gtrsim 0.5^{\circ}$. First of all, we find an enhancement in the nonlinear interaction at the small-$\theta$ regime indicating the strong influence of the band hybridization due to the Umklapp processes. We find strong attractive nonlinearity for hIX while the nonlinear interaction for hX remains repulsive. Intriguingly, these results correlate with measurements in MoS2 bilayers, where redshifts were observed [30]. The emergence of attractive nonlinearity may be understood as a manifestation of strong anisotropic nature of 2D bilayers. This leads to a weaker interlayer e-h interaction as compared to the intralayer e-e and h-h interaction (Fig. 2b, inset). Hence, the total nonlinearity from each of these interacting channels does not cancel leading to stronger attractive nonlinearity. This is different from the monolayer exciton where large cancellations [61, 62, 63] between these channels result in a weaker repulsive nonlinearity ($\sim 0.5$ nm2eV). This anisotropic electronic property is unique for 2D materials that enable the realization of attractive interaction between dipolar excitons. We also note that the strong attractive X-X interaction between hIX in the small $\theta$ regime comes from the exchange scattering involving the excitonic modes in Zone 2 (Fig. 2a, inset). For instance, in the scattering processes with $\bm{g}\neq\tilde{\bm{g}}$ (Fig. 1d), the second term in the last line in Eq.(9) (all particle exchange process, repulsive) is strongly suppressed due to large momentum transfer ($\mathbf{q}=\bm{g}-\tilde{\bm{g}}$), while the exchange interaction $V_{\nu\tilde{\nu}}^{\nu^{\prime}\tilde{\nu}^{\prime}}$ and $V_{\nu\tilde{\nu}}^{\tilde{\nu}^{\prime}\nu^{\prime}}$ remain large. This makes a striking difference between moiré bilayer and untwisted bilayer ($\theta=0$ and $60^{\circ}$) where the scattering processes only take place within Zone 1 with $\bm{g}=\tilde{\bm{g}}=0$. Furthermore, in our calculation, we find weak repulsive direct interaction for hIX in a homobilayer with $\theta\approx 0$ (see SM[52]). This may be expected since a large electrical dipole moment is forbidden due to the approximate inversion symmetry. As a result, the dominant contribution to $g_{X}$ in Fig. 2b comes from the (attractive) exchange interactions. However, this can change for heterobilayers or electrically-biased samples.
@@ Appendix C Momentum integration in direct and exchange interactions
$M^{\vec{r},\vec{t}}_{\vec{m}}(\vec{\lambda})=\Big{[}\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}\frac{(-1)^{r_{i}}m_{i}!(2\lambda_{i})^{m_{i}-2r_{i}}}{r_{i}!t_{i}!(m_{i}-2r_{i}-t_{i})!}\Big{(}\frac{2}{\vec{\lambda}^{2}}\Big{)}^{(m_{i}-2r_{i}-t_{i}+1)/2}\Big{]}\frac{1}{2}\Gamma\Big{(}\frac{\sum_{i=1}^{4}(m_{i}-2r_{i}-t_{i})+1}{2}\Big{)}$ (38)
@@ Appendix D MoS2 homobilayer with $\theta\approx 0$
In this section, we calculate the spectrum of the MoS2 homobilayer with $\theta\approx 0$ in Fig.4. In contrast to the $\theta\approx 60^{\circ}$ case, the attractive interaction in this stacking is stronger because the inversion symmetry forbid large electric dipole moment in a hybridized exciton.
Figure 4: (a) Moiré exciton spectrum with $\theta\approx 0$. (b) the nonlinearity of moiré exciton.